OpenStudy (anonymous):

Simplify: (5.4 x 103)(7.2 x 10-6) 3.888 x 1017 3.888 x 10-3 3.888 x 10-2 3.888 x 1010

OpenStudy (diyadiya):

\[(5.4 \times 10^3)(7.2 \times 10^{-6})\]\[5.4 \times 7.2 \times 10^{-6} \times 10^3\]

OpenStudy (diyadiya):

First can you do 5.4 x 7.2 ?

OpenStudy (anonymous):

38.88

OpenStudy (diyadiya):

Yes ,That's right Now what is \[10^3\times 10^{-6} = ?\]

mathslover (mathslover):

\[(5.4*10^3)(7.2*10^{-6})\] \[54/10 * 10^3)(72/10*/10^6)\] \[5400*7200000\] \[38880000000\] convert it in scientific form , well diya is going on a good way i didn't read the question properly

OpenStudy (anonymous):

9994

OpenStudy (diyadiya):

ok you have to remember this \[\Large x^m \times x^n= x^{m+n}\] I'll give you an example \[10^2 + 10^4 = 10^{2+4} =10^6\] so what is \[\large 10^3 \times 10^{-6} = ?\]

OpenStudy (anonymous):

IM SORRY ITS -3

OpenStudy (diyadiya):

Ok so what is \[\Large 10^{-3} \times 10 ^ {-6}= ?\]

OpenStudy (anonymous):

3.888 X10-3

OpenStudy (diyadiya):

\[ \large(5.4 \times 10^{3})( 7.2 \times 10^{-6}) \] is this your question or \[ \large (5.4 \times 10^{-3})( 7.2 \times 10^{-6}) \]

OpenStudy (anonymous):

THE FIRST ONE

OpenStudy (diyadiya):

Ok so we got \[\large 38.88 \times 10^{3} \times 10^{-6}\]\[\large 10^{3} \times 10^{-6} = 10^{3+(-6)}\] so \[38.88 \times 10^{-3}\]

OpenStudy (diyadiya):

So now we need this in the form 3.888 x

OpenStudy (diyadiya):

so inorder to shift the decimal to the left side we multiply by \(10^{-1}\) or just divide by 10